Bert is twice as Chang.five yrs ago the sum of their ages was 23.how old is bert
Answer by Cromlix(4381)

Hi there,
Bert is twice as Chang
B = 2C
five yrs ago the sum of their ages was 23
B - 5 + C - 5 = 23
Replace B with 2C from first equation.
2C - 5 + C - 5 = 23
Collect like terms
3C - 10 = 23
3C = 23 + 10
3C = 33
C = 11
Bert = 22 yrs old
Chang = 11 yrs old
Hope this helps:-)